A seven-digit number (i.e. a whole number between 1000000 and 9999999) is selected at random.
What is the probability that it contains no digits other than 6's, 7's, and/or 8's?

What is the probability it contains two 6's, three 7's, and two 8's?

Enter your answers as fractions in lowest terms.
help please.

Answer by Edwin McCravy(20054)

First we count the number of 7 digit integers that contain no other
digits than 6's, 7's, or 8's.

Choose the 1st digit 3 ways, either 6, 7, or 8.
Choose the 2nd digit 3 ways, either 6, 7, or 8.
Choose the 3rd digit 3 ways, either 6, 7, or 8.
Choose the 4th digit 3 ways, either 6, 7, or 8.
Choose the 5th digit 3 ways, either 6, 7, or 8.
Choose the 6th digit 3 ways, either 6, 7, or 8.
Choose the 7th digit 3 ways, either 6, 7, or 8. 

That's 3∙3∙3∙3∙3∙3∙3 = 37 = 2187 ways

That's the numerator of the desired probability,
the number of ways we can succeed in drawing a
7 digit integers that contains no other digits 
than 6's, 7's, or 8's.

Now we count the number of ways to choose ANY
7-digit number between 1000000 and 9999999:

Two ways to do that.  We could do it as above, but
the easy way is to realize there are 9999999 integers'
from 1 through 9999999, inclusive, and we eliminate
those between 1 through 999999, those with fewer
than 7 digits, and there are 999999 of those, so we
subtract:

9999999
-999999
-------
9000000 = the denominator of the desired probability,
the number of ways to succeed or fail.

Desired probability = 2187/9000000 which reduces by
dividing top and bottom by 9:

Answer = 243/1000000

-------------------------
 
What is the probability it contains two 6's, three 7's, and two 8's?

There are 7 positions a digit in a 7-digit integer which a digit can
occupy.

There are '7 positions choose 2' or 7C2=21 ways to fill in the two 6's.
There then remain 5 positions to fill. 
There are '5 positions choose 3' or 5C3=10 ways to fill in the three 7's.
There remain 2 positions to fill.
There are '2 positions choose 2' or 2C2=1 way to fill in the two 8's.

Answer = (7C2)(5C3)(2C2) = 21∙10∙1 = 210 ways

Notice it would not have mattered if we had filled in the 7's first:

There are '7 positions choose 3' or 7C3=35 ways to fill in the three 7's.
There then remain 4 positions to fill. 
There are '4 positions choose 2' or 4C2=6 ways to fill in the two 6's.
There remain 2 positions to fill.
There are '2 positions choose 2' or 2C2=1 way to fill in the two 8's.

Answer = (7C3)(4C2)(2C2) = 35∙6∙1 = 210 ways

That was not a necessary step, but just to show you that you can choose
them in any order.

Desired probability = 210/9000000 which reduces by
dividing top and bottom by 30:

Answer = 7/300000

Edwin